OpenStudy (anonymous):

Find and equation for the plane that contains the line x = -1+3t, y = 5+2t, z = 2-t and is perpendicular to the plane 2x-4y+2z=9 Explain your reasonings behind the solution, thanks!

OpenStudy (amistre64):

you are given 2 vectors; one of them is in the line equation itself as the coeffs of the variable t. <3,2,-1> the other vector in the plane is the normal attached to the plane equation that is "perp" to this one; <2,-4,2> we are given a point to work off of from the line equation as well: the constants: (-1,5,2)

OpenStudy (amistre64):

we can cross the vectors to get a normal to our plane and then dot it to our point

OpenStudy (amistre64):

<3,2,-1> <2,-4,2> --------- nx = 2(2)-(-4)(-1) = 0 ny = 2(-1)-3(2) = -8 nz = 3(-4) - 2(2) = -16 <0,-8,-16> looks to be our normal; or <0,8,16> 0(x-Px)+8(y-Py)+16(z-Pz) = 0 8(y-5)+16(z-2) = 0 8y-40+16z-32 = 0 7y+16z-72 = 0 maybe

OpenStudy (amistre64):

7y meant to be 8y

OpenStudy (anonymous):

Thank you, great answer!

OpenStudy (amistre64):

youre welcome, i hope its right; its been awhile since i did these

OpenStudy (anonymous):

It's correct, the only tricky part really, was to realise you can actually use the normal from the plane to cross with the vector from the line. It was there somewhere, but I obviously didnt remember it :D

OpenStudy (amistre64):

:) yay!!